Suppose we dissolve 12.5 g of chiral compound in .2L of solution and place it in a tube that is 20cm long. What is the samples specific rotation of the observed rotation experimentally determined to be 4 degrees?
A) 20 degrees
B) 12.5 degrees
C) .125 degrees
D) 8 degrees
E) 32 degrees
Answer #1

specific rotation \alpha   = observed rotation/l*conc (g/ml)

                         l           = 20cm = 2dm

                         conc   = 12.5/200 = 0.0625g/ml

                                     = 4/2*0.0625

                                       = 4/0.125   = 320

             E) 32 degrees